What is the wavelength of an electron moving at 5.31×10⁶m/sec?


1
Expert's answer
2020-07-14T09:06:44-0400

Given mass of electron is "9.11\times10^{31}kg"

"h=6.626\times10^{34}J.s"

de Broglie's Equation is

"\lambda=\frac{h}{mv}"

"\lambda=\frac{6.626\times10^{34}J.s}{9.11\times10^{31} kg\times5.31\times106m/sec}"


"\lambda=\frac{6.626\times10^{34}J.s}{4.84\times10^{24}kg.m/sec}"

"\lambda=1.37\times10^{10}m"

Therefore the wavelength an electron moving is "1.37\times10^{10}m"